Find the formula for the inverse function for f(x)=7x^3+5
I'm not sure if I did this right but I got 3√y-5/7=x
If not then what steps should I take to get the inverse?
Thank You

Answer by opeodu(24)

Let f(x)=y y=7x^3+5 y-5=7x^3 ([y-5]/7)=x^3 x=([y-5]/7)^1/3 f^-1(x)=([x-5]/7)^1/3